Stackhouse Computing produces high performance desktop computers. Labor cost data shows that the company spent $280,000 for 1,000 computers produced, and each computer requires 2 technician hours and 5 hours of direct labor. Direct labor is paid $14 per hour by the company. What is the cost of 1 technician hour?
Answer #1

Cost per computer = 280,000/1000

= 280 per computer

280 = (2*Technician cost per hour) + (5*14)

2*Technician cost per hour = 280 - 70 = 210

Technician cost per hour = 210/2

= 105